An atom can be described as roughly having the shape of a sphere. The radius of a potassium atom is 280pm. If the potassium atoms could be lined up in a row, how many of them would it take to make a 1.00 ft row of potassium atoms?


1
Expert's answer
2021-01-28T05:35:57-0500

Radius, r = 280 pm

Diameter = 2r 

         = 2×280 pm

         = 560 pm

 1 pm = 10^-12 m

 So, diameter = 560×10^-12 m


Now,

1.00 ft = 0.305 m

So, number of potassium atoms fits with in 1.00 ft = 0.305/(560×10^-12) 

             = 5.45×10^-8 


So, it takes 5.45×10^-8 potassium atoms to make a 1.00 ft row .